# Harmonizing Generalization and Specialization: Uncertainty-Informed Collaborative Learning for Semi-supervised Medical Image Segmentation

### v1

@@ Abstract
Vision foundation models have demonstrated strong generalization in medical image segmentation by leveraging large-scale, heterogeneous pretraining. However, they often struggle to generalize to specialized clinical tasks under limited annotations or rare pathological variations, due to a mismatch between general priors and task-specific requirements. To address this, we propose Uncertainty-informed Collaborative Learning (UnCoL), a dual-teacher framework that harmonizes generalization and specialization in semi-supervised medical image segmentation. Specifically, UnCoL distills both visual and semantic representations from a frozen foundation model to transfer general knowledge, while concurrently maintaining a progressively adapting teacher to capture fine-grained and task-specific representations. To balance guidance from both teachers, pseudo-label learning in UnCoL is adaptively regulated by predictive uncertainty, which selectively suppresses unreliable supervision and stabilizes learning in ambiguous regions. Experiments on diverse 2D and 3D segmentation benchmarks show that UnCoL consistently outperforms state-of-the-art semi-supervised methods and foundation model baselines. Moreover, our model delivers near fully supervised performance with markedly reduced annotation requirements.
@@ 1 Introduction
In parallel, semi-supervised learning (SSL) has been extensively studied to alleviate annotation demands by leveraging unlabeled data, employing techniques such as consistency regularization [11, 12], self-training [13, 14], and uncertainty estimation [4, 15]. However, these methods are typically developed independently of foundation models, limiting their capacity to incorporate external knowledge priors. Recent SAM-enhanced SSL methods [16, 17] attempt to bridge this gap by guiding foundation model predictions through handcrafted or model-generated prompts. However, they often assume that foundation models can accurately capture task-specific semantics. This assumption may not hold under domain shifts or complex anatomical variations, which can lead to unreliable predictions and subsequently inaccurate pseudo-labels for SSL in the absence of explicit reliability assessment.
@@
We conduct comprehensive experiments on diverse 2D and 3D medical segmentation benchmarks, demonstrating that UnCoL consistently achieves state-of-the-art performance under limited annotation settings.
@@ 2.1 Foundation Models in Medical Image Segmentation
Vision foundation models such as SAM [5], MedSAM [6], and SAM-Med3D [7] demonstrate promising generalization in medical image segmentation by leveraging large-scale, heterogeneous pretraining across diverse modalities and anatomical structures. Through sparse prompts (e.g., points or boxes), these models enable zero-shot or few-shot segmentation without task-specific retraining. To improve adaptability to downstream tasks, a variety of adaptation strategies have been explored, including full or partial fine-tuning [9], architectural modifications [18], and parameter-efficient methods such as adapter modules and LoRA [10]. However, these methods often remain prompt-dependent and fail to deliver task-specific precision under domain shifts or limited annotations. This underscores the need for frameworks that can leverage foundation model priors more autonomously and with greater label efficiency.
@@ 2.2 Semi-supervised Medical Image Segmentation
Semi-supervised learning (SSL) is widely adopted in medical image segmentation to alleviate annotation burdens by leveraging unlabeled data alongside limited labeled samples. Representative strategies include consistency regularization [19, 14] and pseudo-labeling [12, 20], with mean teacher framework [3] serving as a cornerstone method. Subsequent works have extended this paradigm by incorporating geometric transformations [21], prototype-based representations [11], and cross-subnet consistency [13]. To improve pseudo-label reliability, uncertainty estimation has been introduced to filter out low-confidence predictions. Techniques such as Monte Carlo dropout [22], model ensembles [14], and entropy-based filtering [4, 15] have shown effectiveness in suppressing noise from unlabeled supervision.
@@
In contrast, UnCoL introduces an uncertainty-informed collaboration that adaptively weights teacher reliability per pixel. Furthermore, rather than aligning probabilistic outputs as in SemiSAM+ [23], UnCoL conducts dual-path representation distillation, transferring visual and semantic features from the frozen foundation encoder to a lightweight student. This design captures transferable priors while enabling prompt-free inference and encourages a natural balance between foundation-level generality and task-specific specialization.
@@ 3.2.1 Model Architecture
Generalized teacher: A frozen foundation model pretrained on large-scale and diverse medical datasets. It provides stable and high-capacity guidance via visual and semantic representations.
@@ 3.3 Dual-Path Knowledge Distillation
We introduce a Dual-Path Knowledge Distillation (DPKD) module to transfer multi-level visual representations and prompt-informed semantics from the frozen foundation model to the student. DPKD consists of two paths: (1) visual distillation from intermediate transformer layers, and (2) semantic distillation via prompt-conditioned features derived from the teacher’s image-prompt feature fusion module, which is used only during the distillation stage. This design enables the student to inherit both visual priors and prompt-informed semantics without requiring prompts during inference. The architecture of the DPKD module is illustrated in Fig. 3. Let and denote the encoders of the teacher and student models, consisting of and layers, respectively. A set of layer pairs is selected for alignment, where a mapping function assigns each selected student layer to a corresponding teacher layer . This multi-layer mapping allows the lightweight student to inherit multi-level representations from the deeper teacher, ensuring structural consistency despite the difference in network depth.
@@ 3.4 Uncertainty-Aware Pseudo-Labeling
Given an unlabeled image defined over a spatial domain , the generalized and specialized teachers independently produce class probability maps , where denotes the number of semantic classes. Shannon entropy is used as an uncertainty measure. For each spatial location , we compute the predictive uncertainty using Shannon entropy. Specifically, the uncertainty from the generalized teacher is given by , and from the specialized teacher as . A binary confidence mask is constructed for each teacher by thresholding the entropy with a ramp-up threshold (see Sec. 4.1), yielding and , where is the indicator function and increases over training (see Sec. 4.1 for details).
@@ 3.5.1 Stage 1: Data-Efficient Pretraining
During this stage, a knowledge adapter is applied to fuse the student’s image features with prompt-informed representations from the foundation model. The adapter is implemented as a cross-attention module, where the student’s image features act as the query and the teacher-provided features as key and value. This design enables semantic conditioning while preserving prompt-free inference.
@@ 3.5.2 Stage 2: Semi-Supervised Fine-Tuning
In the second stage, both labeled and unlabeled data () are used. To exploit spatial structures across domains, we adopt a copy-paste strategy that mixes patches from labeled images with ground-truth annotations and unlabeled images with pseudo-labels, ensuring that the label maps are mixed consistently with the image content. A specialized teacher is constructed by applying an EMA to the student model’s parameters. During both training stages, its parameters are updated by , where and denote the parameters of the specialized teacher and student models, respectively. The momentum coefficient is set to , following [3]. The student continues to be supervised by on labeled samples, in line with [20], and is supervised on unlabeled samples using pseudo-labels generated via the uncertainty-aware strategy described in Sec. 3.4. To preserve generalization, the visual distillation loss is reserved, while semantic distillation is omitted to avoid potential noise from synthetic prompts.

### v2

@@ Abstract
Vision foundation models have demonstrated strong generalization in medical image segmentation by leveraging large-scale, heterogeneous pretraining. However, they often struggle to generalize to specialized clinical tasks under limited annotations or rare pathological variations, due to a mismatch between general priors and task-specific requirements. To address this, we propose Uncertainty-informed Collaborative Learning (UnCoL), a dual-teacher framework that harmonizes generalization and specialization in semi-supervised medical image segmentation. Specifically, UnCoL distills both visual and semantic representations from a frozen foundation model to transfer general knowledge, while concurrently maintaining a progressively adapting teacher to capture fine-grained and task-specific representations. To balance guidance from both teachers, pseudo-label learning in UnCoL is adaptively regulated by predictive uncertainty, which selectively suppresses unreliable supervision and stabilizes learning in ambiguous regions. Experiments on diverse 2D and 3D benchmarks show that UnCoL consistently outperforms existing methods across most datasets and metrics, while achieving comparable performance in only a few cases. Moreover, our model delivers near fully supervised performance with markedly reduced annotation requirements. Code is available at: https://github.com/VivienLu/UnCoL.
@@ 1 Introduction
In parallel, semi-supervised learning (SSL) has been extensively studied to alleviate annotation demands by leveraging unlabeled data, employing techniques such as consistency regularization [39, 7], self-training [3, 32], and uncertainty estimation [41, 22]. However, these methods are typically developed independently of foundation models, limiting their capacity to incorporate external knowledge priors. Recent SAM-enhanced SSL methods [28, 45, 20] attempt to bridge this gap by guiding foundation model predictions through handcrafted or model-generated prompts. However, they often assume that foundation models can accurately capture task-specific semantics. This assumption may not hold under domain shifts or complex anatomical variations, which can lead to unreliable predictions and subsequently inaccurate pseudo-labels for SSL in the absence of explicit reliability assessment.
@@
We conduct comprehensive experiments on diverse 2D and 3D medical segmentation benchmarks, demonstrating that UnCoL achieves strong and competitive performance under limited annotation settings.
@@ 2.1 Foundation Models in Medical Image Segmentation
Vision foundation models such as SAM [18], MedSAM [23], and SAM-Med3D [34] demonstrate promising generalization in medical image segmentation by leveraging large-scale, heterogeneous pretraining across diverse modalities and anatomical structures. Through sparse prompts (e.g., points or boxes), these models enable zero-shot or few-shot segmentation without task-specific retraining. To improve adaptability to downstream tasks, a variety of adaptation strategies have been explored, including full or partial fine-tuning [2], architectural modifications [5], and parameter-efficient methods [43]. However, these methods often remain prompt-dependent and fail to deliver task-specific precision under domain shifts or limited annotations. This underscores the need for frameworks that can leverage foundation model priors more autonomously and with greater label efficiency.
@@ 2.2 Semi-supervised Medical Image Segmentation
Semi-supervised learning (SSL) is widely adopted in medical image segmentation to alleviate annotation burdens by leveraging unlabeled data alongside limited labeled samples. Representative strategies include consistency regularization [27, 32] and pseudo-labeling [7, 38], with mean teacher framework [33] serving as a cornerstone method. Subsequent works have extended this paradigm by incorporating geometric transformations [19], prototype-based representations [39], [21, 35], and cross-subnet consistency [3]. To improve pseudo-label reliability, uncertainty estimation has been commonly introduced to filter low-confidence predictions, where Monte Carlo dropout [37], model ensembles [32], and entropy-based filtering [41, 22] effectively suppress noise from unlabeled supervision. Despite their effectiveness, these approaches are predominantly domain-specific and do not explicitly leverage external foundation-model priors.
@@
In contrast, UnCoL introduces an uncertainty-informed collaboration that adaptively weights teacher reliability per pixel. Furthermore, rather than aligning probabilistic outputs, UnCoL transfers visual and semantic representations from the frozen foundation encoder to a lightweight student. This design captures transferable priors while enabling prompt-free inference and encourages a natural balance between foundation-level generality and task-specific specialization.
@@ 3.3 Dual-Path Knowledge Distillation
We introduce a Dual-Path Knowledge Distillation (DPKD) module to transfer multi-level visual representations and prompt-informed semantics from the frozen foundation model to the student. DPKD includes two paths: (1) visual distillation from intermediate transformer layers, and (2) semantic distillation via prompt-conditioned features derived from the teacher’s image-prompt feature fusion module used only during pretraining. This design enables the student to inherit visual priors and prompt-informed semantics without requiring prompts during inference. DPKD architecture is illustrated in Fig. 3. Let and denote the encoders of the teacher and student models, consisting of and layers, respectively. A set of layer pairs is selected for alignment, where a mapping function assigns each selected student layer to a corresponding teacher layer . This multi-layer mapping allows the lightweight student to inherit multi-level representations from the deeper teacher, ensuring consistency despite differences in network depth.
@@ 3.4 Uncertainty-Aware Pseudo-Labeling
Given an unlabeled image defined on a spatial domain , the generalized and specialized teachers independently produce class probability maps , where denotes the number of semantic classes. Predictive uncertainty at each location is computed using Shannon entropy. Specifically, the uncertainties from the generalized and specialized teachers are given by and , respectively. A binary confidence mask is constructed for each teacher by thresholding entropy with a ramp-up threshold , yielding and , where is the indicator function and increases over training (see Sec. 4.1 for details).
@@ 3.5.1 Stage 1: Data-Efficient Pretraining
During this stage, a knowledge adapter is applied to fuse the student’s image features with prompt-informed representations from the foundation model. The adapter is implemented as a cross-attention module, where student’s image features act as the query and teacher-provided features as key and value. This design enables semantic conditioning while preserving prompt-free inference.
@@ 3.5.2 Stage 2: Semi-Supervised Fine-Tuning
In the second stage, both labeled and unlabeled data () are used. To exploit spatial structures across domains, we adopt a copy-paste strategy that mixes patches from labeled images with ground-truth annotations and unlabeled images with pseudo-labels, ensuring that the label maps are mixed consistently with the image content. A specialized teacher is constructed by applying EMA to the student model’s parameters. During both training stages, its parameters are updated by , where and denote the parameters of the specialized teacher and student models, respectively. The momentum coefficient is set to , following [33]. The student continues to be supervised by on labeled data, in line with [38], and is supervised on unlabeled samples using pseudo-labels generated via the uncertainty-aware strategy. To preserve generalization, the visual distillation loss is reserved, while semantic distillation is omitted to avoid potential noise from synthetic prompts.